Homogenous Linear Differential Equations

6. (D ^ 3 - 2D ^ 2 - 3D) * y = 0

5. (D ^ 3 - 3D ^ 2 + 4) * y = 0

7. (4D ^ 3 - 3D + 1) * y = 0


1
Expert's answer
2022-02-15T15:28:01-0500

D3 - 2D2 - 3D"\times"y = 0


(D-1)(D2-2D)=0


D=1, "\frac{2+\sqrt(4-8)}{2}"


y(x)= "\frac{1}{D^3-3D^2+4D-2}" ex


y(x)= "\frac{1}{1-3+4-2}" ex


y= "\frac{1}{1-3+4-2}" xex



D3 - 3D2 + 4)"\times" y = 0


(D-1)(D2-3D+4)=0


D=1, "\frac{2+\sqrt(4-8)}{2}"


y(x)= "\frac{1}{D^3-3D^2+4D-2}" ex


y(x)= "\frac{1}{1-3+4-2}" ex


y= "\frac{1}{1-3+4-2}" xex



(4D3 - 3D + 1)"\times" y = 0


(D-1)(4D2-3D+1)=0


D=1, "\frac{3+\sqrt(4-3)}{2}"


y(x)= "\frac{1}{D^3-3D^2+4D-2}" ex


y(x)= "\frac{1}{1-3+4-2}" ex


y= "\frac{1}{1-3+4-2}" xex